OpenStudy (anonymous):

If you know 40% of a number, explain how you could work out the original number.

OpenStudy (anonymous):

Let the number n be 40% => n/40 = 1% => 100n/40 = 100%....the original amount. In this case this becomes 2.5n.............as 2.5 x 40 = 100

OpenStudy (anonymous):

thanks for the answer

OpenStudy (anonymous):

yw:) thanks for the medal.

OpenStudy (anonymous):

your welcome :)

OpenStudy (anonymous):

let b be 40% of the number X. \[\frac{ 40 }{ 100 }timesX=b\] \[0.4timesX=b\] \[X =\frac{ b }{0.4 }\]